You've lost a lawsuit that requires you to pay $1,500 per month for
120 months starting next month.
How much would you have to invest today to exactly pay for the lawsuit, if your investments earn 3.00% APR (compounded monthly)?

Calculate your answer to the nearest dollar.
Answer #1

This question requires application of PV of annuity, according to which:

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

r = 3%/12 = 0.25% (monthly), n = 120, P = $1500

PV = 1500-1-(1 + 0.0025)-120 0.0025

PV = 1500 * 103.5618

PV = $155,343